# Gaussian Herding across Pens:::: An Optimal Transport Perspective on Global Gaussian Reduction for 3DGS

### v1

@@ Appendix D Additional Scene Visualizations
In addition to the previously shown Figure 4, we present more detailed qualitative comparisons across multiple scenes in Figure 8. Our method consistently preserves the visual quality of the original models. Notably, when applied to stronger 3DGS variants with improved densification strategies—such as Mini-Splatting-D—our compaction framework performs even better. This is reflected in the fact that, after compaction, Mini-Splatting-D often achieves higher rendering quality than the original 3DGS baseline.
Table 4: Additional information in plug-in experiment.
Method-Compact Rate $\rho$ Tank&Temples MipNeRF-360 Deep Blendering
SSIM PSNR LPIPS $k$ Gaussians FPS SSIM PSNR LPIPS $k$ Gaussians FPS SSIM PSNR LPIPS $k$ Gaussians FPS
3DGS-30k 10% 0.818 23.312 0.242 157 341 0.764 26.404 0.314 272 272 0.905 29.647 0.264 248 290
20% 0.835 23.615 0.212 313 300 0.788 26.973 0.275 545 230 0.907 29.864 0.252 496 237
100% 0.853 23.785 0.169 1577 168 0.813 27.554 0.221 2625 126 0.907 29.816 0.238 2476 126
Mini-Splatting-D 10% 0.835 23.232 0.198 147 564 0.802 27.090 0.25 318 389 0.909 30.042 0.254 156 549
20% 0.855 23.403 0.171 293 526 0.821 27.310 0.214 636 299 0.912 30.17 0.238 318 489
100% 0.848 23.338 0.14 1473 209 0.832 27.486 0.176 3177 111 0.907 29.980 0.211 1597 197
Atom-GS 10% 0.793 22.988 0.274 426 401 0.764 26.535 0.307 468 304 0.899 29.347 0.282 464 419
20% 0.812 23.282 0.24 852 283 0.788 27.025 0.269 937 236 0.902 29.347 0.268 925 323
100% 0.814 23.289 0.235 4274 87 0.796 27.135 0.251 4699 86 0.902 29.314 0.267 4625 121

### v2

@@ Appendix E Additional Numerical Results and Scene Visualizations
In addition to the previously shown Figure 5, we present more detailed qualitative comparisons across multiple scenes in Figure 8. Our method consistently preserves the visual quality of the original models. In particular, when applied to stronger 3DGS variants with improved densification strategies, such as Mini-Splatting-D, our compaction framework performs even better. This is reflected in the fact that, after compaction, Mini-Splatting-D often achieves higher rendering quality than the original 3DGS baseline.
Table 4: We compared GHAP with four post-processing methods (LightGaussian, PUP-3DGS, Trimming the Fat, MesonGS) at 20% retention, as well as two end-to-end methods (LocoGS, 3DGS-MCMC). GHAP also replaces the pruning in Mini-Splatting. Results show that our method substantially outperforms both post-processing and end-to-end approaches.
Method Tanks&Temples MipNeRF-360 Deep Blending
[…]
original 3DGS 0.853 23.785 0.169 1577 0.813 27.554 0.221 2627 0.907 29.816 0.238 2475
LocoGS 0.843 23.655 0.191 571 0.798 27.049 0.257 674 0.903 29.972 0.261 529
3DGS+GHAP (ours) 0.835 23.615 0.212 314 0.788 26.973 0.275 527 0.907 29.864 0.252 496
LightGaussian-20% 0.779 22.486 0.271 315 0.765 26.353 0.288 526 0.873 28.011 0.315 495
PUP-3DGS-20% 0.809 22.603 0.228 315 0.790 26.671 0.257 525 0.905 29.719 0.248 495
Trimming the Fat-20% 0.819 22.498 0.232 315 0.781 26.494 0.280 524 0.900 29.082 0.272 494
MesonGS-20% 0.822 20.699 0.207 314 0.776 25.006 0.262 527 0.897 28.696 0.262 496
3DGS-MCMC 0.779 22.141 0.282 315 0.763 25.957 0.309 263 0.885 28.976 0.298 496
MiniSplatting-20% 0.824 22.953 0.223 142 0.794 26.728 0.267 215 0.904 29.763 0.265 240
MiniSplatting+GHAP (ours) 0.855 23.403 0.171 155 0.821 27.310 0.214 219 0.912 30.170 0.238 250